OpenStudy (anonymous):

4_01]Which of the following is the slope between the points (0, -7) and (4, 12)? 19/4 -19/4 4/19 -4/19

OpenStudy (apoorvk):

\[\text{The slope 'm' between two points A(}x_1,y_1\text{) and B(}x_2,y_2\text{) is:}\] \[m=\frac{y_2 - y_1}{x_2 - x_1}\] Whaddya think now?

OpenStudy (anonymous):

slope is y/x y=12-(-7)=19 x=4-0=4 so, it is 19/4

OpenStudy (anonymous):

thankz